OpenStudy (anonymous):

who knows 2 derieve the equation of motion"v^2=u^2+2as

OpenStudy (anonymous):

i bet there arent many!!!!

OpenStudy (anonymous):

u guys fail 2 me hahhaha

OpenStudy (anonymous):

s=vt v=u+v/2 s=((u+v)/2)/t--1 we know that v=u+at--2 plugging that in equation 1 s=[(u+u+at)/2]/t s=[(2u+at)/2]/t s=ut+1/2at^2--3 from equation 2 v^2=(u+at)^2 v^2=u^2+2uat+(at)^2 v^2=u^2+2a(ut+(1/2)at^2) using equation 3 we have v&2=u^2+2as

OpenStudy (anonymous):

hreat which grade r u???????

OpenStudy (anonymous):

v^2=u^2+2aS*

OpenStudy (anonymous):

i know its physics

OpenStudy (anonymous):

Yh it is physics:)

OpenStudy (anonymous):

r u 12th????

OpenStudy (anonymous):

or college

OpenStudy (anonymous):

i am in biomedical engineering

OpenStudy (anonymous):

actually going for it

OpenStudy (anonymous):

oh tats y i know it now itself!!!!!! derievation i am just 6th

OpenStudy (anonymous):

good:)

OpenStudy (anonymous):

what a humble kid

OpenStudy (jamesj):

Alternatively, by the Work-Energy theorem) (change in energy) = (work) = (Force).(displacement force applied) \[\frac{1}{2} m v^2 - \frac{1}{2} m u^2 = F s = mas\] Canceling m throughout and multiplying by 2 \[v^2 - u^2 = 2as\]